The mass of the planet Jupiter is approximately 1.8981 X 10^27 kilograms. The mass of the planet Venus is approximately 4.8673 x 10^24 kilograms. About how many times larger is the mass of Jupiter than that of Venus?
Found 2 solutions by MathLover1, Alan3354:
Answer by MathLover1(20849)


Jupiter is approximately 1.8981* 10^27 kilograms. The mass of the planet Venus is approximately 4.8673* 10^24 kilogram
Jupiter/Venus=1.8981* 10^27/4.8673* 10^24
Jupiter/Venus=(1.8981/4.8673) 10^(27-24)
Jupiter/Venus=0.38996979845* 10^3
Jupiter/Venus=389.96979845
Jupiter/Venus=389.97
the mass of the planet Jupiter is 389.97 larger than the mass of the planet Venus


Answer by Alan3354(69443)

The mass of the planet Jupiter is approximately 1.8981 X 10^27 kilograms. The mass of the planet Venus is approximately 4.8673 x 10^24 kilograms. About how many times larger is the mass of Jupiter than that of Venus?
-------------
Divide
1.8981e27/4.8672e24 = ~ 0.38998e3 = 389.98 times AS LARGE, not larger
= 388.98 times as large.
============================
Times larger is not the same as "times as large."
1 time larger = twice as large, an increase of 100%
----
When the numbers a big, the difference between, eg, 397 and 398 is less significant.